OpenStudy (anonymous):

Simplify the expression the root of negative one all over the quantity of four minus three i minus the quantity of two plus five i. .

OpenStudy (anonymous):

sqrt(-1)/(4-3i)-(2+5i)

OpenStudy (anonymous):

this is impossible since you have a negative inside the root

OpenStudy (anonymous):

square root of a -x is equal to xi

OpenStudy (anonymous):

first symplify the fraction i/(4-3i)=(i(4+3i))/((4+3i)(4-3i)

OpenStudy (anonymous):

(4i+3i^2)/(16-9i^2)

OpenStudy (anonymous):

\[\sqrt{-4}\] this equals to 2i for example \[\sqrt{4}\times \sqrt{-1}\] \[\sqrt{4}=2\] \[\sqrt{-1}=i\]

OpenStudy (anonymous):

(4i-3)/(16+9)=(4i-3)/25

OpenStudy (anonymous):

I would send the choices but .. they are too confusing because everything is typed out.